OpenStudy (anonymous):

Use scientific notation to perform the calculations. (8.4 x 10^-13) (4.8 x 10^9) I got one answer, but the book got another and I am not sure how I messed it up. Help?

OpenStudy (anonymous):

what dd u get

OpenStudy (anonymous):

I got 4.03 x 10^-4 and the book said it was 4.032 x 10^-3

OpenStudy (anonymous):

(8.4 x 10^-13) (4.8 x 10^9) = (8.4 x 4.8)(10^-13 x 10^9) = (40.32)(10^-13+9) = 40.32 x 10^-4 but in scientific notation, we leave only one number to the left of the decimal point, so the decimal has to be shifted left to achieve this we do the following shown in attached draw file

OpenStudy (anonymous):

press ur calculator very well, i also got \[4.032\times10^{-3}\]

OpenStudy (anonymous):